OpenStudy (anonymous):

Simplify the expression: \[\frac{ 8x ^{-4}y ^{-8} }{ -2xy ^{5} }\] Write your answer without negative exponents.

OpenStudy (anonymous):

I know I just asked a similar question. Right now I am not clear on how to handle it if the denominator is not the same as the numerator as demonstrated above.

sam (.sam.):

You gotta understand your indices well, recall \[\Large \frac{a^{m}}{a^{n}}=a^{m-n}\]

OpenStudy (anonymous):

Okay, so how would I apply that to this question? I saw that in the lesson and tried to understand, but I really can't see how it applies here.

sam (.sam.):

So this \[\frac{ 8x ^{-4}y ^{-8} }{ -2xy ^{5} }\] Is simply \[\Large -4(x^{-4-1} )( y^{-8-5})\]

sam (.sam.):

There's a few way but this is a snappier one

OpenStudy (anonymous):

So \[-4 \left(x ^{3}\right)\left( y ^{3} \right)\]

sam (.sam.):

\[-4(x^{-4-1} )( y^{-8-5}) \\ \\ =-4(x^{-5} )( y^{-13})\] Then use \[\Large x^{-a}=\frac{1}{x^{a}}\] \[\Large -4(\frac{1}{x^5y^{13}})\]

sam (.sam.):

-4-1 is -5 and -8-5 is -13

OpenStudy (anonymous):

\[-4\left( x ^{5}y ^{13} \right)\]?

sam (.sam.):

\[\LARGE -4(x^{-4-1} )( y^{-8-5}) \\ \\ \LARGE =-4(x^{-5} )( y^{-13})\]

sam (.sam.):

\[\LARGE =-4(x^{-5} y^{-13})\]

sam (.sam.):

Then use \[\Large x^{-a}=\frac{1}{x^{a}}\]

sam (.sam.):

You'll get \[\Large -4(\frac{1}{x^5y^{13}})\]

OpenStudy (anonymous):

And thats the answer?

sam (.sam.):

Yep

OpenStudy (anonymous):

Ah, okay... thanks!